# hyperrefSuppressing link with empty target

### v1

@@ 3.3 Horizontal-Wind and Temperature Profiles
Moving deeper into the atmosphere( 5(d)) we start to find evidence that horizontal advection is shaping the thermal structure as the influence of radiation wanes. Of particular note are the equatorial winds, which are now easterly at all longitudes, although the influence of divergent day-night winds remains with the equatorial winds being significantly slower just west of the substellar point (see subsection 3.4 for more details). We also catch our first glimpse of a rather unusual feature that is exclusive to the slowly rotating regime: a slight hot-spot on the night-side, where the divergent winds converge (as we discuss in Figure 9, this is not a coincidence).
@@
This night-side hot-spot grows in strength as we move towards mid-pressures ($0.2$\mathrm{bar}$$ 5(g)), resulting in a rather unusual wind and temperature structure: the hottest region of the atmosphere has moved from the day-side to just west of the anti-stellar point, where the divergent day-night winds converge. Further, this night-side hot-spot, and the resulting temperature gradients, drive a significant equatorial wind from the night-side to the day-side, which appears in the zonal-mean zonal-wind profile as part of the jet, despite it not forming in the same way as the classical hot Jupiter zonal-mean equatorial jet.
@@
If robust, the formation of this night-side hot spot, which occurs for all of our HD209458b-like models that fall into the slowly rotating regime, may have important implications for our understanding of the dynamics of such objects. For example, as discussed in subsection 3.5 and shown in 10(a), the night-side hot-spot is associated with a thermal inversion, the presence of which might have strong implications for both the observed phase curve (depending upon the wavelength/molecule observed, and hence the pressure probed) as well as the chemistry, since the inversion can act as a cold-trap, leading to condensation/rain-out and hence changing the, observable, composition of the atmosphere (which in turn may effect the thermal structure due to changes in absorption/emission). Note that, as discussed above, this feature is common to all our slowly rotating regime models, and, in fact, its presence is almost a defining feature of the slowly rotating regime. This is because, as $\Omega$ increases, rotational influences shape the day-night flow, shifting the hot-spot even further westwards, away from the anti-stellar point and towards the terminator, until it eventually combines with the easterly advected day-side hot-spot, leading to the formation of the thermal butterfly that is emblematic of the classical hot-Jupiter regime at a transitional rotation rate of around $\Omega=0.5\rightarrow 0.6\Omega_{0}$.
@@
For example, in the outer atmosphere (5(b)), we find slight advection from the day-side towards the eastern terminator, forming ‘wings’ that: a) trace the off-equator divergent/convergent flows, which are increasingly rotationally influenced, and hence have become longitudinally asymmetric, with eastern flows compressed and, higher latitude, western flows stretched longitudinally; and b) will eventually form, deeper in the atmosphere, a key part of the thermal butterfly structure that is synonymous with the classical regime. We also find that, unlike in the slowly rotating regime, super-rotating flows are present at all longitudes on the equator, albeit with the proviso that winds west of the sub-stellar point are slightly suppressed/weakened by the divergent component of the wind (see subsection 3.4).
Evidence for the formation of foreshadowed thermal butterfly increases as we move deeper, 5(e), with the influence that rotation has on the dynamics also becoming increasingly clear. For example, the high latitude flows have become increasingly asymmetric, shifting the ‘wings’ associated with convergent flows closer towards the terminator and the day-side hot-spot. Together wind the growing strength of the equatorial jet at all longitudes (although it does remain slightly weaker just west of the sub-stellar point) which drives the eastward advection (by approximately $20^{\circ}$) of the near equator day-side hot-spot, we are starting to see the basis of the structure that will form the thermal butterfly.
A structure that becomes apparent as we move into the mid atmosphere ($0.2$\mathrm{bar}$$ 5(h)), where we find that the day-side hot-spot is now located $\sim 50^{\circ}$ east of the sub-stellar point. Interestingly, we also find a secondary peak in the temperature profile even further east, across the terminator and on the night-side - this is a remnant of the ‘wings’ seen at lower pressures and can also be considered as a sign that the off-equator heat transport was sufficiently rotationally influenced as to prevent the formation of the night-side hot-spot found in the slowly rotating regime. Further, whilst the off-equator flows are mostly hidden by the equatorial jet, the visible remnants suggest that, at high latitudes, the easterly circulation that previously had been confined to the day-side has been almost completely suppressed, leaving only a westerly day-night flow, which appears in the zonal-mean zonal-wind as a counterflow (1(b)).
@@
Here, in the outer atmosphere (5(c)) we find that the off-equator wind dynamics are very strongly suppressed (by the Coriolis effect), particularly on the cold night-side, resulting in only a weak day-side circulation from west of the sub-stellar point to the east via mid-latitudes. This leaves the strong super-rotating equatorial jet as the dominant flow, but even this is influenced by rotation, with the jet being visibly narrower than its classical regime counterpart. Consequently, this highly rotationally-confined wind structure drives thermal advection that is limited to equatorial regions, with weak eastwards advection at the equator (weak due to the reduced jet speed that latitudinally confined standing waves can drive), weak westwards advection at low latitudes, and basically no heat transport at high latitudes (i.e. above ±$\pm\sim 30^{\circ}$).
@@ 3.4 Helmholtz Wind Decomposition
where $\bm{k}$ denotes the unit vector in the (eastward) zonal direction, $\chi$ is the velocity potential function, $\psi$ is the velocity stream-function, and both $\chi$ and $\psi$ can be linked to the divergence $\delta$ / vorticity $w$ directly:
@@
We start by analysing the winds in the slowly rotating regime, which is plotted in the left column of Figure 6. Here, our analysis reveals that, in the outer and mid atmosphere, the primary component which make up the total horizontal wind is the divergent component ($\left|\bm{u}_{d}\right|=554\mathrm{\,ms^{-1}}$), which is, on average, a factor of two faster than the rotational component ($\left|\bm{u}_{r}\right|=291\mathrm{\,ms^{-1}}$). This makes sense, and is in good agreement with the wind structure discussed in subsection 3.1. To start, off-equator, we find flows that trace the dominant, divergent, day-side to night-side flow (compare 6(a) and 5(a)), and which are driven by the strong radiative forcing and day-night temperature gradient. Whereas at the equator, we find that the zonal wind associated with rotational dynamics opposes the divergent wind west of the sub-stellar point whilst also reinforcing the flow to the east, leading to to observed asymmetry in equatorial wind speed. Decomposing the rotational wind further reveals the driving force behind this equatorial jet: as shown in 6(g), the eddy component of the wind consists of an $m=1$ circulation structure (i.e. two circulation cells per hemisphere), a structure which is associated with Kelvin and Rossby standing waves and which is commonly attributed as being the driving mechanism behind super-rotating jets on hot Jupiters (Showman & Polvani 2011). Briefly, this mechanism works as follows: longitudinal variations in the radiative forcing (i.e. the strong day/night heating/temperature contrast) trigger the formation of standing, planetary-scale, Rossby and Kelvin waves, with the Kelvin waves straddling the equator, and the Rossby waves lying at mid latitudes (with their exact location depending upon the rotation rate). This standing wave pattern then essentially acts to pump eastward angular momentum from mid-latitudes to the equator, inducing equatorial superrotation and mid-latitude counterflows. Note that this angular momentum pumping is also associated with an eastwards tilt of the standing wave pattern, a tilt which is strongly associated with Coriolis forces, and hence is only found at more rapid rotation rates (6(h)-l).
@@
As we increase the rotation rate and move into the classical hot Jupiter regime, shown in the middle column of Figure 6, we find that the balance has shifted in favour of the rotational component ($\left|\bm{u}_{r}\right|=692\mathrm{\,ms^{-1}}$) over the divergent component ($\left|\bm{u}_{d}\right|=241\mathrm{\,ms^{-1}}$), with the eddy component (6(h) - $\left|\bm{u}_{e}\right|=283\mathrm{\,ms^{-1}}$) starting to exhibit signs of the titled standing wave pattern associated with a strong rotationally influenced zonal jet, as seen in 6(e) and subsection 3.1. This dominance of rotation driven dynamics is also reflected in the thermal/wind structures shown in Figure 5. For example, comparing 6(e) with 5(b)/f/g reveals the thermal butterfly formation mechanism: a strong jet drives equatorial advection whilst the $m=1$ circulations drives the formation of the off-equator temperature wings east of the sub-stellar point as well as the slightly cooler point found on the equator just behind the tip of the butterfly. Both of these features are associated with the rotational wind, with the divergent wind (6(b)) only acting to slightly slowdown/speedup the jet west/east of the sub-stellar point.
@@
Moving into the rapidly rotating regime, as shown on right-hand column of Figure 6, we start to see strong evidence of the Coriolis driven reshaping of the winds. For example, we find that the rotational wind ($\left|\bm{u}_{r}\right|=455\mathrm{\,ms^{-1}}$) completely dominates the divergent wind ($\left|\bm{u}_{d}\right|=32\mathrm{\,ms^{-1}}$), leading to a divergent wind whose contributions are insignificant at all but the lowest of pressures where radiative forcing dominates. However this change has not come about because of an enhancement of the rotational wind (which is slower here than either slower rotation regime), rather we find that Coriolis effects have started to seriously suppress off-equator flows, essentially destroying the divergent wind and leaving a highly latitudinally compressed and extremely titled eddy wind (6(i) - $\left|\bm{u}_{e}\right|=83\mathrm{\,ms^{-1}}$) that struggles to generate a broad zonal jet, but does still manage to drive at least some off-equator advection east of the sub-stellar point (e.g. 5(c)). Evidence for this compression can be seen in 1(c), where we find that the off-equator counter flows associated with the standing wave pattern have moved to low latitudes. We also find that, off-equator, both components of the wind are notably weaker on the night-side than the day-side, an effect which we attribute to the tilt of the eddy winds and the standing wave pattern leading to a preferred direction for circulations and essentially zero night-side flow.
@@
It is worth mentioning that the compression/suppression of the off-equator flows only grows in strength as we increase the rotation rate further. This has devastating consequences for overall zonal advection/transport: as the standing-wave structure gets further compressed/suppressed by Coriolis forces, its ability to pump off-axis, eastwards, angular momentum towards the equator is reduced, resulting in a slight slowdown in equatorial jet speeds, as shown in 7(a), as well as a narrowing of the jet, as shown in 2(b). Together this leads to a significant reduction is zonal transport except at the equator, the effects of which can be seen in Figure 5. As we discuss below, this can have significant consequences for the vertical advection of enthalpy in rapidly rotating atmospheres, and may help to explain the lack of inflation observed in such objects.
(a) $R_{0}$ vs. $U\right>_{equator}$
(b) $R_{0}$ vs. $W\right>_{pole}$
Figure 8: Global mean global vertical enthalpy flux for 29 HD209458b-like models spanning the rotation range $0.125\Omega_{0}\rightarrow 40\Omega_{0}$. Note that negative fluxes indicate net downward enthalpy (potential temperature) transport.
@@
(d) $0.125\Omega_{0}$ Vertical - 0.016 bar
(e) $1\Omega_{0}$ Zonal - 0.0026 bar
(f) $1\Omega_{0}$ Zonal - 0.1 bar
(g) $1\Omega_{0}$ Meridional - 0.016 bar
(h) $1\Omega_{0}$ Vertical - 0.016 bar
(i) $15\Omega_{0}$ Zonal - 0.0026 bar
(j) $15\Omega_{0}$ Zonal - 0.1 bar
(k) $15\Omega_{0}$ Meridional - 0.016 bar
(l) $15\Omega_{0}$ Vertical - 0.016 bar
@@
Figure 10: Latitudinally averaged (over the equatorial regions) T-P profiles for three HD209458b-like models at different rotation rates; $0.125\Omega_{0}$ top, $1\Omega_{0}$ middle, and $15\Omega_{0}$ bottom. Each plot includes profiles from 6 different longitudes, ranging from the sub-stellar point (whose equilibrium profile is shown in red) eastwards to the anti-stellar point (whose equilibrium profile is shown in blue) on the night-side. Note that we have excluded the deepest regions of the atmosphere ($P>50\textrm{bar}$) from these plots so as to emphasise the changes that occur with longitude at lower pressures, whilst still showing how the deep atmospheric heating changes (compared to the initial adiabat, shown as a dark grey dash) with rotation rate.
@@
Figure 11: Longitudinal variations (top) and global means (bottom) of the vertical enthalpy flux $F_{\mathcal{H}}$ profiles for three exemplary HD209458b-like models; $0.125\Omega_{0}$ top, $\Omega_{0}$ middle, and $4\Omega_{0}$ bottom. In the top row, we plot vertical enthalpy flux profiles at 5 different longitudes, ranging from the sub-stellar point to just west of the anti-stellar point, as well as the mean enthalpy flux profile. However, since the mean flux is significantly smaller than the local fluxes, we replot the mean profiles in the bottom row such as to better demonstrate the vertical variations in enthalpy transport.

### v2

@@ 3.3 Horizontal-Wind and Temperature Maps
Moving deeper into the atmosphere( 5(d)) we start to find evidence that horizontal advection is shaping the thermal structure as the influence of radiation wanes. Of particular note are the equatorial winds, which are now eastwards at all longitudes, although the influence of divergent day-night winds remains with the equatorial winds being significantly slower just west of the substellar point (see subsection 3.4 for more details). We also catch our first glimpse of a rather unusual feature that is exclusive to the slowly rotating regime: a slight hot-spot on the night-side, where the divergent winds converge (as we discuss in Figure 7, this is not a coincidence).
@@
This night-side hot-spot grows in strength as we move towards mid-pressures ($0.2$\mathrm{bar}$$ 5(g)), resulting in a rather unusual wind and temperature structure: the hottest region of the atmosphere has moved from the day-side to just west of the anti-stellar point, where the divergent day-night winds converge. Further, this night-side hot-spot, and the resulting temperature gradients, drive a significant equatorial wind from the night-side to the day-side, which appears in the zonal-mean zonal-wind profile as part of the jet, despite it not forming in the same way as the classical hot Jupiter zonal-mean equatorial jet. In part, these differences may occur due to differences in the behaviour of Rossby waves, and barotropic and baroclinic instabilities, as the Rossby deformation radius approaches the planetary radius (Mitchell & Vallis 2010).
@@
If robust, the formation of this night-side hot spot, which occurs for all of our HD209458b-like models that fall into the slowly rotating regime, may have important implications for our understanding of the dynamics of such objects. For example, as discussed in subsection 3.5 and shown in 10(a), the night-side hot-spot is associated with a thermal inversion, the presence of which might have strong implications for both the observed phase curve (depending upon the wavelength/molecule observed, and hence the pressure probed) as well as the chemistry, since the inversion can act as a cold-trap, leading to condensation/rain-out and hence changing the, observable, composition of the atmosphere (which in turn may effect the thermal structure due to changes in absorption/emission). Note that, as discussed above, this feature is common to all our slowly rotating regime models, and, in fact, its presence is almost a defining feature of the slowly rotating regime. This is because, as $\Omega$ increases, rotational influences shape the day-night flow, shifting the hot-spot even further westwards, away from the anti-stellar point and towards the terminator, until it eventually combines with the eastwards advected day-side hot-spot, leading to the formation of the thermal butterfly that is emblematic of the classical hot-Jupiter regime at a transitional rotation rate of around $\Omega=0.5\rightarrow 0.6\Omega_{0}$. This shift in the location of the hot-spot is in agreement with the Spitzer phase curve observations analysed by May et al. (2022), which revealed (as shown in their Figure 10) an increasing trend of hot-spot off set with increasing orbital period (i.e. decreasing synchronous rotation rate). They even found a slight westward shift at the most rapid rotation rates, a shift which might correspond to the off-equator westwards advection that we find in the rapidly rotating regime (see below).
@@
For example, in the outer atmosphere (5(b)), we find slight advection from the day-side towards the eastern terminator, forming ‘wings’ that: a) trace the off-equator divergent/convergent flows, which are increasingly rotationally influenced, and hence have become longitudinally asymmetric, with eastern flows compressed and, higher latitude, western flows stretched longitudinally; and b) will eventually form, deeper in the atmosphere, a key part of the thermal butterfly structure that is synonymous with the classical regime. Note that this structure is also known as a global ‘chevron’ pattern (Showman & Polvani 2011). We also find that, unlike in the slowly rotating regime, super-rotating flows are present at all longitudes on the equator, albeit with the proviso that winds west of the sub-stellar point are slightly suppressed/weakened by the divergent component of the wind (see subsection 3.4).
Evidence for the formation of foreshadowed thermal butterfly increases as we move deeper, 5(e), with the influence that rotation has on the dynamics also becoming increasingly clear. For example, the high latitude flows have become increasingly asymmetric, shifting the ‘wings’ associated with convergent flows closer towards the terminator and the day-side hot-spot. Together with the growing strength of the equatorial jet at all longitudes (although it does remain slightly weaker just west of the sub-stellar point) which drives the eastward advection (by approximately $20^{\circ}$) of the near equator day-side hot-spot, we are starting to see the basis of the structure that will form the thermal butterfly.
A structure that becomes apparent as we move into the mid atmosphere ($0.2$\mathrm{bar}$$ 5(h)), where we find that the day-side hot-spot is now located $\sim 50^{\circ}$ east of the sub-stellar point. Interestingly, we also find a secondary peak in the temperature even further east, across the terminator and on the night-side - this is a remnant of the ‘wings’ seen at lower pressures and can also be considered as a sign that the off-equator heat transport was sufficiently rotationally influenced as to prevent the formation of the night-side hot-spot found in the slowly rotating regime. Further, whilst the off-equator flows are mostly hidden by the equatorial jet, the visible remnants suggest that, at high latitudes, the easterly circulation that previously had been confined to the day-side has been almost completely suppressed, leaving only a westerly day-night flow, which appears in the zonal-mean zonal-wind as a counterflow (1(b)).
@@
Here, in the outer atmosphere (5(c)) we find that the off-equator wind dynamics are very strongly suppressed (by the Coriolis effect), particularly on the cold night-side, resulting in only a weak day-side circulation from west of the sub-stellar point to the east via mid-latitudes. This leaves the strong super-rotating equatorial jet as the dominant flow, but even this is influenced by rotation, with the jet being visibly narrower than its classical regime counterpart. Consequently, this highly rotationally-confined wind structure drives thermal advection that is limited to equatorial regions, with weak eastwards advection at the equator (weak due to the reduced jet speed that latitudinally confined standing waves can drive), weak westwards advection at low latitudes, and basically no heat transport at high latitudes (i.e. above ±$\pm\sim 30^{\circ}$). This weak low-latitude westwards advection may explain the slight westward phase curve offset found for short orbital period hot Jupiters by May et al. (2022).
@@ 3.4 Helmholtz Wind Decomposition
where $\bm{k}$ denotes the unit vector in the (eastward) zonal direction, $\chi$ is the velocity potential function, $\psi$ is the velocity stream-function, and both $\chi$ and $\psi$ can be linked to the divergence $\delta$ / vorticity $\mathcal{W}$ directly:
@@
We start by analysing the winds in the slowly rotating regime, which is plotted in the left column of Figure 6. Here, our analysis reveals that, in the outer and mid atmosphere, the primary component which make up the total horizontal wind is the divergent component ($\left|\bm{u}_{d}\right|=554\mathrm{\,ms^{-1}}$), which is, on average, a factor of two faster than the rotational component ($\left|\bm{u}_{r}\right|=279\mathrm{\,ms^{-1}}$). This makes sense, and is in good agreement with the wind structure discussed in subsection 3.1. To start, off-equator, we find flows that trace the dominant, divergent, day-side to night-side flow (compare 6(a) and 5(a)), and which are driven by the strong radiative forcing and day-night temperature gradient. Whereas at the equator, we find that the zonal wind associated with rotational dynamics opposes the divergent wind west of the sub-stellar point whilst also reinforcing the flow to the east, leading to to observed asymmetry in equatorial wind speed. Decomposing the rotational wind further reveals the driving force behind this equatorial jet: as shown in 6(g), the eddy component of the wind consists of an $m=1$ circulation structure (i.e. two circulation cells per hemisphere), a structure which is associated with Kelvin and Rossby standing waves and which is commonly attributed as being the driving mechanism behind super-rotating jets on hot Jupiters (Showman & Polvani 2011). Briefly, this mechanism works as follows: longitudinal variations in the radiative forcing (i.e. the strong day/night heating/temperature contrast) trigger the formation of standing, planetary-scale, Rossby and Kelvin waves, with the Kelvin waves straddling the equator, and the Rossby waves lying at mid latitudes (with their exact location depending upon the rotation rate). Rotational effects (i.e. Coriolis forces) then cause this standing wave pattern to tilt, with a northwest-southwest tilt in the northern hemisphere and a southwest-northeast tilt in the southern hemisphere. These tilted standing waves pump eastwards angular momentum from mid-latitudes to the equator, inducing equatorial superrotation and mid-latitude counterflows. Note that thus tilt is essential to this angular momentum transporting mechanism and hence an equatorial jet is typically only found at more rapid rotation rates (6(h)-l).
@@
As we increase the rotation rate and move into the classical hot Jupiter regime, shown in the middle column of Figure 6, we find that the balance has shifted in favour of the rotational component ($\left|\bm{u}_{r}\right|=692\mathrm{\,ms^{-1}}$, where $\left|\bm{x}\right|$ indicates the RMS mean of $\bm{x}$) over the divergent component ($\left|\bm{u}_{d}\right|=241\mathrm{\,ms^{-1}}$), with the eddy component (6(h) - $\left|\bm{u}_{e}\right|=283\mathrm{\,ms^{-1}}$) starting to exhibit signs of the titled standing wave pattern associated with a strong rotationally influenced zonal jet, as seen in 6(e) and subsection 3.1. This dominance of rotation driven dynamics is also reflected in the thermal/wind structures shown in Figure 5. For example, comparing 6(e) with 5(b)/f/g reveals the thermal butterfly formation mechanism: a strong jet drives equatorial advection whilst the $m=1$ circulations drives the formation of the off-equator temperature wings east of the sub-stellar point as well as the slightly cooler point found on the equator just behind the tip of the butterfly. Both of these features are associated with the rotational wind, with the divergent wind (6(b)) only acting to slightly slowdown/speedup the jet west/east of the sub-stellar point.
@@
Moving into the rapidly rotating regime, as shown on right-hand column of Figure 6, we start to see strong evidence of the Coriolis driven reshaping of the winds. For example, we find that the rotational wind ($\left|\bm{u}_{r}\right|=377\mathrm{\,ms^{-1}}$) completely dominates the divergent wind ($\left|\bm{u}_{d}\right|=40\mathrm{\,ms^{-1}}$), leading to a divergent wind whose contributions are insignificant at all but the lowest of pressures where radiative forcing dominates. However this change has not come about because of an enhancement of the rotational wind (which is slower here than either slower rotation regime), rather we find that Coriolis effects have started to seriously suppress off-equator flows, essentially destroying the divergent wind and leaving a highly latitudinally compressed and extremely titled eddy wind (6(i) - $\left|\bm{u}_{e}\right|=98\mathrm{\,ms^{-1}}$) that struggles to generate a broad zonal jet, but does still manage to drive at least some off-equator advection east of the sub-stellar point (e.g. 5(c)). Evidence for this compression can be seen in 1(c), where we find that the off-equator counter flows associated with the standing wave pattern have moved to low latitudes. We also find that, off-equator, both components of the wind are notably weaker on the night-side than the day-side, an effect which we attribute to the tilt of the eddy winds and the standing wave pattern leading to a preferred direction for circulations and essentially zero night-side flow.
@@
It is worth mentioning that the compression/suppression of the off-equator flows only grows in strength as we increase the rotation rate further. This has devastating consequences for overall zonal advection/transport: as the standing-wave structure gets further compressed/suppressed by Coriolis forces, its ability to pump off-axis, eastwards, angular momentum towards the equator is reduced, resulting in a slight slowdown in equatorial jet speeds, as shown in 17(a), as well as a narrowing of the jet, as shown in 2(b). Together this leads to a significant reduction is zonal transport except at the equator, the effects of which can be seen in Figure 5. As we discuss below, this can have significant consequences for the vertical advection of enthalpy in rapidly rotating atmospheres, and may help to explain the lack of inflation observed in such objects.
@@
(d) $1\Omega_{0}$ Zonal - 0.0026 bar
(e) $1\Omega_{0}$ Zonal - 0.1 bar
(f) $1\Omega_{0}$ Meridional - 0.016 bar
(g) $15\Omega_{0}$ Zonal - 0.0026 bar
(h) $15\Omega_{0}$ Zonal - 0.1 bar
(i) $15\Omega_{0}$ Meridional - 0.016 bar
@@
(a) $0.125\Omega_{0}$ Vertical - 0.016 bar
(b) $0.125\Omega_{0}$ Vertical - 0.1 bar
(c) $0.125\Omega_{0}$ Vertical - 4 bar
(d) $1\Omega_{0}$ Vertical - 0.016 bar
(e) $1\Omega_{0}$ Vertical - 0.1 bar
(f) $1\Omega_{0}$ Vertical - 4 bar
(g) $15\Omega_{0}$ Vertical - 0.016 bar
(h) $15\Omega_{0}$ Vertical - 0.1 bar
(i) $15\Omega_{0}$ Vertical - 4 bar
Figure 8: Continued from previous page: Maps showing the Zonal, Meridional, and Vertical Enthalpy transport at select pressure levels for three exemplary HD209458b-like models with rotation rates of; $0.125\Omega_{0}$ top, $\Omega_{0}$ middle, and $15\Omega_{0}$ bottom. Here, positive (red) fluxes represent eastward/polar/outwards flows for the zonal/meridional/vertical enthalpy flux maps respectively. Note: We include two maps of the zonal enthalpy transport at different pressure levels in order to emphasise how the zonal advection changes with height.
(a) $0.125\Omega_{0}$ Vertical Wind
(b) $0.125\Omega_{0}$ Temperature
(c) $0.125\Omega_{0}$ Vertical Enthalpy
(d) $1\Omega_{0}$ Vertical Wind
(e) $1\Omega_{0}$ Temperature
(f) $1\Omega_{0}$ Vertical Enthalpy
(g) $15\Omega_{0}$ Vertical Wind
(h) $15\Omega_{0}$ Temperature
(i) $15\Omega_{0}$ Vertical Enthalpy
Figure 9: Maps showing longitudinal slices of the Vertical Wind (left), Temperature (middle), and Vertical Enthalpy Flux (right), at the equator, for three exemplary HD209458b-like models with rotation rates of; $0.125\Omega_{0}$ top, $\Omega_{0}$ middle, and $15\Omega_{0}$ bottom. Here we have confined our plots to pressures less than 0.26 bar in order to emphasise the dynamics and transport in the outer and mid atmosphere, particularly at pressure levels similar to those at which the night-side hot-spot forms. Note the difference in scale for the vertical enthalpy flux in our $15\Omega_{0}$ case, emphasising the weaker vertical transport, even at the equator, in the rapidly rotating regime.
@@
Figure 10: Latitudinally averaged ( over a latitudinal band spanning ±$\pm 10^{\circ}$ of the equator) T-P profiles for three HD209458b-like models at different rotation rates; $0.125\Omega_{0}$ top, $1\Omega_{0}$ middle, and $15\Omega_{0}$ bottom. Each plot includes profiles from 6 different longitudes, ranging from the sub-stellar point (whose equilibrium profile is shown in red) eastwards to the anti-stellar point (whose equilibrium profile is shown in blue) on the night-side. Note that we have excluded the deepest regions of the atmosphere ($P>50\textrm{bar}$) from these plots so as to emphasise the changes that occur with longitude at lower pressures, whilst still showing how the deep atmospheric heating changes (compared to the initial adiabat, shown as a dark grey dash) with rotation rate.
@@
Figure 11: Longitudinal variations (top) and global means (bottom) of the vertical enthalpy flux $F_{\mathcal{H}}$ profiles for three exemplary HD209458b-like models; $0.125\Omega_{0}$ left, $\Omega_{0}$ middle, and $15\Omega_{0}$ right. In the top row, we plot vertical enthalpy flux profiles at 5 different longitudes, ranging from the sub-stellar point to just west of the anti-stellar point, as well as the mean enthalpy flux profile. However, since the mean flux is significantly smaller than the local fluxes, we replot the mean profiles in the bottom row such as to better demonstrate the vertical variations in enthalpy transport.